OpenStudy (britnee_paige99):

F(x)=3x+5/x what is f(a+2)

OpenStudy (mathmale):

You are to "evaluate F(x) at the x value a-2. 1. Write F(x) = 3x + 5/x as given. 2. if you wish, re-write this by eliminating each incidence of 'x' and replacing it with an empty set of parentheses: F( ) = ? Please share your results. 3. Now fill in the empty sets of parentheses with 'a-2' 4. Simplify the right side of your equation. Done.

OpenStudy (calculusxy):

\[\large \frac{ 3(a+2) + 5 }{ a + 2 } = \frac{ 3a + 6 + 5 }{ a + 2 } = \frac{ 3a + 11 }{ a + 2 }\]